# Generalized Schalkwijk-Kailath Coding for Autoregressive Gaussian Channels

### v1

@@ 1 Introduction and Summary
    We propose the SK(2) coding scheme for AR($p$) Gaussian channels, some type of generalized SK coding, and derive its achievable rate (Theorem 4.1, Corollary 4.2).
@@
    For AR(2) Gaussian channels, the SK(2) coding scheme can strictly outperform the celebrated SK coding scheme, thereby disproving a conjecture of Butman [2] (Example 4.6).
@@ 2 Notation and Terminologies
We use $\delta_{ij}$ to denote the Kronecker delta function. For $a\in\mathbb{C}$, $|a|$ denotes its modulus; for $\in\mathbb{C}^{n}, a\|a\|$ denotes its $\ell^{2}$-norm. For an $m\times n$ matrix $A$, we use $A^{\ast}$ to denote its conjugate transpose (or $A^{T}$ if $A$ is real). If $A$ is square, $|A|$ denotes its determinant, with the convention that $|A|=1$ for an empty matrix (i.e., $m=n=0$).
@@ 3 The Schalkwijk–Kailath Coding
where $\{V_{i}\}$ follows the deterministic recursion
@@
Let $\{u_{k};k=1,2,...,m\}\subset\mathbb{C}^{n}$ for $n,m\geq 1$. Then, it holds that
$u_{i}^{\ast}(I_{n}+UU^{\ast})^{-1}u_{j}=\left\{\begin{aligned} 1\penalty 10000\ -\penalty 10000\ \frac{M_{ii}}{|A|}&\qquad\mbox{if $i=j$};\\ (-1)^{i+j+1}\frac{M_{ji}}{|A|}&\qquad\mbox{if $i\neq j$},\end{aligned}\right$.
where $U=\begin{bmatrix}u_{1}&u_{2}&...&u_{m}\end{bmatrix}$ and $A=I_{m}+U^{\ast}U$ and $M_{ij}$ is the $(i,j)$-minor of $A$, i.e., the determinant of the $(m-1)\times(m-1)$ matrix that results from deleting row $i$ and column $j$ of $A$. In particular, if $m=1$, then we have
$u_{1}^{\ast}(I_{n}+UU^{\ast})^{-1}u_{1}=1-1/A.$
The following theorem provides the maximal achievable rate $\bar{I}_{\mathrm{SK1}}(P)$ for the SK(1) coding scheme over stationary AR(p)(p) Gaussian channels, yielding a lower bound on $C_{\mathrm{SK1}}(P)$.
@@
Then, the maximal achievable rate for the SK(1) coding scheme $\bar{I}_{\mathrm{SK1}}(P)$ is given by
$\bar{I}_{\mathrm{SK1}}(P)\triangleq\max_{\gamma}\log|\gamma|,$ (12)
where the maximum is taken over all $|\gamma|>1$ satisfying the power constraint
@@ 4 The SK(2) Coding Scheme
We now state the corresponding achievable rate $\bar{I}_{\mathrm{SK2}}(P)$ for the SK(2) coding scheme under the condition $b\neq 0$ in (15) over stationary AR(p)(p) Gaussian channels, yielding a lower bound on $C_{\mathrm{SK2}}(P)$.
@@
Then, under the condition $b\neq 0$ in (15), the maximal achievable rate for the SK(2) coding $\bar{I}_{\mathrm{SK2}}(P)$ is given by
$\bar{I}_{\mathrm{SK2}}(P)\triangleq\max_{\gamma_{1},\gamma_{2}}2\log|\gamma_{1}|\wedge|\gamma_{2}|$ (19)
where the maximum is taken over all $|\gamma_{1}|>1,\penalty 10000\ |\gamma_{2}|>1$ satisfying the power constraint
$\frac{1}{\Delta}\left(\frac{L_{11}}{\gamma^{2}_{2}}+\frac{L_{22}}{\gamma_{1}^{2}}+\frac{2L_{12}}{\gamma_{1}\gamma_{2}}\right)\leq P,$ (20)
@@
$\left\{\begin{aligned} \widetilde{Y}_{n}&=Y_{n}^{\ast},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ n\leq p\\ \widetilde{Y}_{n}&=L(B)Y^{\ast}_{n},\penalty 10000\ \penalty 10000\ \penalty 10000\ n\geq p+1\end{aligned}\right$. (22)
@@
The mutual information between $V_{1}^{n}$ and $Y_{1}^{n}$ for $n\geq p+1$ is given by
@@
$=\frac{1}{n}I(U_{1};Y_{1}^{n})+\frac{1}{n}I(U_{2};Y_{1}^{n}|U_{1})$
$\geq\frac{1}{n}I(U_{1};Y_{1}^{n})+\frac{1}{n}I(U_{2};Y_{1}^{n})$
$=\frac{1}{n}I(U_{1};\widetilde{Y}_{1}^{n})+\frac{1}{n}I(U_{2};\widetilde{Y}_{1}^{n})$
$\geq\frac{1}{n}I(U_{1};\widetilde{Y}_{p+1}^{n})+\frac{1}{n}I(U_{2};\widetilde{Y}_{p+1}^{n})$
$=-\frac{1}{2n}\log\mathbb{E}[|U_{1}-\mathbb{E}[U_{1}|\widetilde{Y}_{p+1}^{n}]|^{2}]$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -\frac{1}{2n}\log\mathbb{E}[|U_{2}-\mathbb{E}[U_{2}|\widetilde{Y}_{p+1}^{n}]|^{2}].$
On the other hand, for $n\geq p+2$,
@@
$=\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{1}^{n-1}]|^{2}]$
$\leq E[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{p+1}^{n-1}]|^{2}].$
Without loss of generality, our analysis focuses only on the limiting behavior of the inequalities (25) and (26); therefore, it suffices to redefine the sequence $\{\widetilde{Y}_{k}\}$ in (22) by the form given in (23), taking the initial index $n=1$ rather than $n=p+1$:
$\widetilde{Y}_{n}=d_{1,n}U_{1}+d_{2,n}U_{2}+W_{n}$
for $n\geq 1$. Since $U_{1}$ and $\{\widetilde{Y}_{1}^{n}\}$ are jointly Gaussian, it holds that
@@
$\mathbb{E}[|U_{2}-\mathbb{E}[U_{2}|\widetilde{Y}_{1}^{n}]|^{2}]=\frac{1+D_{1,n}^{T}D_{1,n}}{|I_{2}+D_{n}^{T}D_{n}|}.$ (29)
@@
As a result, consider the rates of convergence of (28) and (29). It turns out from (31) and (32) and (35) that
[…]
or equivalently,
$\lim_{n\to\infty}-\frac{1}{2n}\log\mathbb{E}\left[\left|U_{1}-\mathbb{E}[U_{1}|\widetilde{Y}_{1}^{n}]\right|^{2}\right]=\log(|\gamma_{1}|\wedge|\gamma_{2}|).$ (36)
Similarly, we also have
$\lim_{n\to\infty}-\frac{1}{2n}\log\mathbb{E}\left[\left|U_{2}-\mathbb{E}[U_{2}|\widetilde{Y}_{1}^{n}]\right|^{2}\right]=\log(|\gamma_{1}|\wedge|\gamma_{2}|).$ (37)
Therefore, combining (25), (36), and (37), it follows that the mutual information rate achieved by the SK(2) coding scheme specified in equations (14) and (15) can be lower bounded by
$\lim_{n\to\infty}\frac{1}{n}I(V_{1}^{n};Y_{1}^{n})\geq 2\log(|\gamma_{1}|\wedge|\gamma_{2}|),$ (38)
which establishes the desired lower bound for the mutual information rate achieved by the SK(2) coding scheme.
@@
$=L_{11}\gamma_{1}^{2}\gamma_{2}^{2(n-1)}(\gamma_{1}^{2(n-1)}-1)+L_{22}\gamma_{2}^{2}\gamma_{1}^{2(n-1)}(\gamma_{2}^{2(n-1)}-1)+2L_{12}\gamma_{1}\gamma_{2}\gamma_{1}^{n-1}\gamma_{2}^{n-1}(\gamma_{1}^{n-1}\gamma_{2}^{n-1}-1).$
@@
$\varlimsup_{n\to\infty}\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{1}^{n-1}]|^{2}] lim$$\leq\varlimsup_{n\to\infty}\frac{a_{n}^{2}+b_{n}^{2}+\|a_{n}D_{2,n-1}-b_{n}D_{2,n-1}\|^{2}}{|I_{2}+D_{n-1}^{T}D_{n-1}|}$
@@
$\leq P,$
where the first inequality follows from (35) and (39). This establishes the power constraint (26) for the SK(2) coding scheme. Together with (38), this completes the proof for Case A.
Next, we proceed to establish the result for Case B. Recall that in this case we have $\gamma_{1}=\gamma_{2}=\gamma\in\mathbb{R}$ with $|\gamma|>1$, which yields $a_{n}=(2-n)\gamma^{n-1}$ and $b_{n}=(n-1)\gamma^{n-2}$ as in (18). In the same way we define $\{\widetilde{Y}\}$ as in (22) and thus for $n\geq p+1$
@@
$d_{2,n}$ $=\left[(n-1)L(\gamma^{-1})-L^{\prime}(\gamma^{-1})\gamma^{-1}\right]\gamma^{n-2}$
for $n\geq 1$. Consequently, the scalar products $D^{T}_{1,n}D_{1,n}$, $D^{T}_{2,n}D_{2,n}$ and $D^{T}_{1,n}D_{2,n}$ are readily computed as follows:
@@
$=\frac{1}{\gamma^{2}-1}n^{2}\gamma^{2n}-\frac{2\gamma^{2}}{(\gamma^{2}-1)^{2}}n\gamma^{2n}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{\gamma^{4}(\gamma^{2}+1)}{(\gamma^{2}-1)^{3}}\gamma^{2(n-1)}-\frac{(\gamma^{2}+1)\gamma^{2}}{(\gamma^{2}-1)^{3}}.$
Proceeding analogously to (34) and (35), we derive
@@
Applying this together with (40) (41) and (42) to both (28) and (29), and subsequently taking the limit as $n\to\infty$ as in (36) and (37), leads to
$\lim_{n\to\infty}-\frac{1}{2n}\log\mathbb{E}[|U_{1}-\mathbb{E}[U_{1}|\widetilde{Y}_{1}^{n}]|^{2}]=\log|\gamma|$
and
$\lim_{n\to\infty}-\frac{1}{2n}\log\mathbb{E}[|U_{2}-\mathbb{E}[U_{2}|\widetilde{Y}_{1}^{n}]|^{2}]=\log|\gamma|,$
which immediately implies that, upon taking the limit $n\to\infty$, inequality (25) reduces to
[…]
This inequality provides the desired lower bound on the mutual information rate achieved by the SK(2) coding scheme in Case B.
@@
$\varlimsup_{n\to\infty}\mathbb{E}[|X_{n}|^{2}] =lim$$=\varlimsup_{n\to\infty}\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{1}^{n-1}]|^{2}]$ (45)
[…]
@@
where the second equality follows from the definitions of $a_{n}$ and $b_{n}$ together with (43) and (44).
@@
This immediately establishes the power constraint (20) for Case B. The proof of the identity (46) is deferred to Appendix C. ∎
Now, combining Theorem 3.2 (the case $b=0$) and Theorem 4.1 (the case $b\neq 0$), we now present the main achievable-rate result of this paper for the SK(2) coding scheme, which provides a lower bound on $C_{\mathrm{SK2}(P)}$.
@@
$C_{\mathrm{SK2}}(P)\geq\max\left\{\bar{I}_{\mathrm{SK1}}(P),\bar{I}_{\mathrm{SK2}}(P)\right\},$
where $\bar{I}_{\mathrm{SK1}}(P)$ and $\bar{I}_{\mathrm{SK2}}(P)$ are given by (12) and (19), respectively.
@@
This spectral density corresponds to a stationary noise process of the form $Z_{i}=W_{i},\penalty 10000\ i\geq 1$, where $\{W_{i}\}_{i=1}^{\infty}$ is a zero-mean, unit-variance white Gaussian process. This corresponds precisely to the AWGN channel (4). It follows directly from Theorem 3.2 and Corollary 4.2 that for the case $b=0$, the SK(2) scheme reduces to the SK(1) scheme and is therefore optimal for this channel, as noted in Remark 3.3. For the case $b\neq 0$, the following corollary demonstrates that, for such channel, the optimal pair $(\gamma_{1},\gamma_{2})$ in the optimization problem (19)-(20) satisfies $|\gamma_{1}|=|\gamma_{2}|$. Consequently, the SK(2) coding scheme yields a broader family of optimal coding scheme for the AWGN channel.
@@
Consider the noise process $\{Z_{i}\}_{i=1}^{\infty}$ with the power spectral density given in (47). Then, the feedback capacity of the AWGN channel $Y_{i}=X_{i}+Z_{i},\penalty 10000\ i=1,2,$..., subject to the power constraint $P$, can be achieved via the SK(2) coding scheme defined in (15) and (14) for $b\neq 0$ with any parameter pair $(\gamma_{1},\gamma_{2})$ satisfying $|\gamma_{1}|=|\gamma_{2}|$.
@@
$\bar{I}_{\mathrm{SK2}}(\gamma_{1},\gamma_{2})\triangleq 2\log|\gamma_{1}|\wedge|\gamma_{2}|$
under the constraint (48). Next, we consider the two cases. First, suppose the characteristic roots $(\gamma_{1},\gamma_{2})$ satisfy $\gamma_{1}=\overline{\gamma_{2}}$ and $|\gamma_{2}|=\gamma>1$. In this case, $\bar{I}_{\mathrm{SK2}}(\gamma_{1},\gamma_{2})=2\log|\gamma|$ and the power constraint (48) reduces to $\gamma^{4}-1\leq P$. Consequently,
@@
$\max_{\begin{subarray}{c}\gamma_{1},\gamma_{2}\in\mathbb{R}\\ \gamma_{1}\neq\gamma_{2}\end{subarray}}\bar{I}_{\mathrm{SK2}}(\gamma_{1},\gamma_{2})$ $=\max 2\log|\gamma_{1}|\wedge|\gamma_{2}|$
@@
where the maximum is attained by any pair $(\gamma_{1},\gamma_{2})$ satisfying $\gamma_{1}=\pm(P+1)^{1/4}$ and $\gamma_{2}=\pm(P+1)^{1/4}$.
In summary, combining both cases we have shown that the maximum $\bar{I}_{\mathrm{SK2}}(P)$ in (19) is achieved whenever the characteristic roots $(\gamma_{1},\gamma_{2})$ satisfy $|\gamma_{1}|=|\gamma_{2}|=(P+1)^{1/4}$. In particular, the well-known feedback capacity of the AWGN channel, $\frac{1}{2}\log(P+1)$, is attained by the SK(2) coding scheme for the case $b\neq 0$ with any parameter pair $(\gamma_{1},\gamma_{2})$ satisfying $|\gamma_{1}|=|\gamma_{2}|$, as desired. ∎
@@
with equality if and only if $\beta_{1}=0$. For different values of $\beta_{1}$ and $P=1,5$, we plot $\bar{I}_{\mathrm{SK1}}(P)$, $\bar{I}_{\mathrm{SK2}}(P)$, and their difference as functions of $\beta_{1}$, as shown in Fig. 1.
Figure 1: Comparison of $\bar{I}_{\mathrm{SK1}}(P)$ and $\bar{I}_{\mathrm{SK2}}(P)$ for stationary AR(1) Gaussian channels with $P=1,5$.
@@
$C_{\mathrm{SK2}}(P)\geq\bar{I}_{\mathrm{SK2}}(P)>\bar{I}_{\mathrm{SK1}}(P).$
Thus, unlike for AR(1) Gaussian channels, the $\mathrm{SK}(1)$ coding scheme is not universally optimal. This disproves the conjecture of Butman [2] that in our notation $\bar{I}_{\mathrm{SK1}}(P)=C_{\mathrm{FB}}(P)$.
Figure 2: Comparison of $\bar{I}_{\mathrm{SK1}}(P)$ and $\bar{I}_{\mathrm{SK2}}(P)$ for stationary AR(2) Gaussian channels with $\beta_{2}=0.5,\penalty 10000\ 0.9$ and $P=1$.
@@
For any nonzero $x\in\mathbb{C}^{n}$, it holds that
$x^{\ast}(I_{n}+UU^{\ast})x$ $=x^{\ast}x+x^{\ast}UU^{\ast}x$
$=\|x\|^{2}+\|U^{\ast}x\|^{2}$
$\geq\|x\|^{2}$

### v2

@@ 1 Introduction and Summary
    We propose the SK(2) coding scheme for AR($p$) Gaussian channels, a second-order generalization of the SK coding, and derive its maximal achievable rate (Theorem 4.1, Corollary 4.2).
@@
    For AR(2) Gaussian channels, the SK(2) coding scheme can strictly outperform the celebrated SK coding scheme, thereby disproving the conjecture of Butman [2] (Example 4.6).
@@ 2 Notation and Terminologies
We use $\delta_{ij}$ to denote the Kronecker delta function. For $a\in\mathbb{C}$, $|a|$ denotes its modulus; for $\in\mathbb{C}^{n}, a\|a\|$ denotes its $\ell^{2}$-norm. For any two real numbers $a$ and $b$, we denote their minimum by $a\wedge b$. For an $m\times n$ matrix $A$, we use $A^{\ast}$ to denote its conjugate transpose (or $A^{T}$ if $A$ is real). If $A$ is square, $|A|$ denotes its determinant, with the convention that $|A|=1$ for an empty matrix (i.e., $m=n=0$).
@@ 3 The Schalkwijk–Kailath Coding
where $\{V_{i}\}$ follows the deterministic recursion:
@@
Let $\{u_{k};k=1,2,...,m\}\subset\mathbb{C}^{n}$ for $n,m\geq 1$ and $S$ be a positive-definite Hermitian matrix with $n\times n$. Then for any $1\leq i,j\leq m$,
$u_{i}^{\ast}(S+UU^{\ast})^{-1}u_{j}=\left\{\begin{aligned} 1\penalty 10000\ -\penalty 10000\ \frac{M_{ii}}{|A|}&\qquad\mbox{if $i=j$};\\ (-1)^{i+j+1}\frac{M_{ji}}{|A|}&\qquad\mbox{if $i\neq j$},\end{aligned}\right$.
where $U=\begin{bmatrix}u_{1}&u_{2}&...&u_{m}\end{bmatrix}$ and $A=I_{m}+U^{\ast}S^{-1}U$ and $M_{ij}$ is the $(i,j)$-minor of $A$. In particular, if $m=1$, then we have
$u_{1}^{\ast}(S+UU^{\ast})^{-1}u_{1}=1-1/A.$
The following theorem gives a closed-form expression for the SK(1) channel capacity $C_{\mathrm{SK1}}(P)$ over stationary AR(p)(p) Gaussian channels, which consequently provides a computable lower bound on the feedback capacity $C_{\mathrm{FB}}(P)$.
@@
Then, we have
$C_{\mathrm{SK1}}(P)=\max_{\gamma}\log|\gamma|,$ (12)
where the maximum is taken over all $\gamma\in\mathbb{R}$ with $|\gamma|>1$ satisfying the power constraint
@@ 4 The SK(2) Coding Scheme
Under the condition $b\neq 0$ in (15), we now state a lower bound $\bar{I}_{\mathrm{SK2}}(P)$ on the SK(2) channel capacity $C_{\mathrm{SK2}}(P)$ of the SK(2) coding over the AR($p$) Gaussian channel (2).
@@
Then, we have
$C_{\mathrm{SK2}}(P)\geq\bar{I}_{\mathrm{SK2}}(P),$
where
$\bar{I}_{\mathrm{SK2}}(P)\triangleq\max_{\gamma_{1},\gamma_{2}}\log|\gamma_{1}\gamma_{2}|$ (19)
and the maximum is taken over all $\gamma_{1},\gamma_{2}\in\mathbb{C}$ with $|\gamma_{1}|\wedge|\gamma_{2}|>1$ satisfying $\gamma_{1}+\gamma_{2}\in\mathbb{R}$, $\gamma_{1}\gamma_{2}\in\mathbb{R}$, and the power constraint
$\frac{1}{\Delta}\left(\frac{L_{11}}{\gamma^{2}_{2}}+\frac{L_{22}}{\gamma_{1}^{2}}+\frac{2L_{12}}{\gamma_{1}\gamma_{2}}\right)=P,$ (20)
@@
$\widetilde{Y}_{n}=\left\{\begin{aligned} &Y_{n}^{\ast},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ n\leq p\\ &L(B)Y^{\ast}_{n},\penalty 10000\ \penalty 10000\ \penalty 10000\ n\geq p+1\end{aligned}\right$. (22)
@@
Note that by (22) and (23), the process $\{\widetilde{Y}_{n}\}$ admits two different expressions on either side of $n=p$. However, it turns out from the proof of Lemma 3.2, the initial part $\widetilde{Y}_{1}^{p}$ does not affect the argument. Hence, without loss of generality, we may extend the definition of $\widetilde{Y}_{n}$ in (23) to all $n\geq 1$ instead of only $n\geq p+1$:
$\widetilde{Y}_{n}=d_{1,n}U_{1}+d_{2,n}U_{2}+W_{n},\penalty 10000\ n\geq 1.$ (25)
The mutual information between $V_{1}^{n}$ and $Y_{1}^{n}$ for $n\geq 2$ is given by
@@
$=\frac{1}{n}I(U_{1},U_{2};\widetilde{Y}_{1}^{n})$
$=\frac{1}{n}I(U_{1};\widetilde{Y}_{1}^{n})+\frac{1}{n}I(U_{2};\widetilde{Y}_{1}^{n}|U_{1})$
$=-\frac{1}{2n}\log\mathbb{E}[|U_{1}-\mathbb{E}[U_{1}|\widetilde{Y}_{1}^{n}]|^{2}]-\frac{1}{2n}\log\mathbb{E}[|U_{2}-\mathbb{E}[U_{2}|U_{1},\widetilde{Y}_{1}^{n}]|^{2}].$
On the other hand, for $n\geq 2$,
@@
$=\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{1}^{n-1}]|^{2}].$
Since $U_{1}$ and $\{\widetilde{Y}_{1}^{n}\}$ are jointly Gaussian, it holds that
@@
$\mathbb{E}[|U_{2}-\mathbb{E}[U_{2}|U_{1},\widetilde{Y}_{1}^{n}]|^{2}]$ $=1-\begin{bmatrix}0&D^{T}_{2,n}\end{bmatrix}\begin{bmatrix}1&D_{1,n}^{T}\\ D_{1,n}&I_{n}+D_{1,n}D_{1,n}^{T}+D_{2,n}D_{2,n}^{T}\end{bmatrix}^{-1}\begin{bmatrix}0\\ D_{2,n}\end{bmatrix}$ (30)
$=1-D_{2,n}^{T}\left(I_{n}+D_{2,n}D_{2,n}^{T}\right)^{-1}D_{2,n}$
$=\frac{1}{|1+D_{2,n}^{T}D_{2,n}|}$
Consequently, it follows from (29) and (30) that (26) can be simplified as
$\frac{1}{n}I(V_{1}^{n};Y_{1}^{n})=\frac{1}{2n}\log|I_{2}+D_{n}^{T}D_{n}|.$ (31)
@@
As a result, it follows from (33), (34) and (37) that, upon taking the limit as $n\to\infty$ in (31), the mutual information rate achieved by the SK(2) coding scheme specified in equations (14) and (15) is given by
$\lim_{n\to\infty}\frac{1}{n}I(V_{1}^{n};Y_{1}^{n})=2\log|\gamma_{1}\gamma_{2}|.$ (38)
@@
$=L_{11}\gamma_{1}^{2}\gamma_{2}^{2(n-1)}(\gamma_{1}^{2(n-1)}-1)+L_{22}\gamma_{2}^{2}\gamma_{1}^{2(n-1)}(\gamma_{2}^{2(n-1)}-1)+2L_{12}\gamma_{1}^{n}\gamma_{2}^{n}(\gamma_{1}^{n-1}\gamma_{2}^{n-1}-1).$
@@
$\lim_{n\to\infty}\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{1}^{n-1}]|^{2}]$ $=\lim_{n\to\infty}\frac{a_{n}^{2}+b_{n}^{2}+\|a_{n}D_{2,n-1}-b_{n}D_{2,n-1}\|^{2}}{|I_{2}+D_{n-1}^{T}D_{n-1}|}$ (40)
@@
$=P,$
where the first inequality follows from (37) and (39). Therefore, under the condition $b\neq 0$, it follows from (38) and (40) that the desired results in (19) and (20) hold, completing the proof for Case A.
Next, we proceed to establish the result for Case B. Recall that in this case we have $\gamma_{1}=\gamma_{2}=\gamma\in\mathbb{R}$ with $|\gamma|>1$, which yields $a_{n}=(2-n)\gamma^{n-1}$ and $b_{n}=(n-1)\gamma^{n-2}$ as in (18). In the same way we define $\{\widetilde{Y}\}$ as in (22) and thus for $n\geq 1$
@@
$d_{2,n}$ $=\left[(n-1)L(\gamma^{-1})-L^{\prime}(\gamma^{-1})\gamma^{-1}\right]\gamma^{n-2}.$
Consequently, the scalar products $D^{T}_{1,n}D_{1,n}$, $D^{T}_{2,n}D_{2,n}$ and $D^{T}_{1,n}D_{2,n}$ are readily computed as follows:
@@
$=\frac{1}{\gamma^{2}-1}n^{2}\gamma^{2n}-\frac{2\gamma^{2}}{(\gamma^{2}-1)^{2}}n\gamma^{2n}+\frac{\gamma^{4}(\gamma^{2}+1)}{(\gamma^{2}-1)^{3}}\gamma^{2(n-1)}-\frac{(\gamma^{2}+1)\gamma^{2}}{(\gamma^{2}-1)^{3}}.$
Proceeding analogously to (36) and (37), it follows from (41), (42) and (43) that
@@
Applying this to (31), and subsequently taking the limit as $n\to\infty$, leads to
$\lim_{n\to\infty}\frac{1}{n}I(V_{1}^{n};Y_{1}^{n})=2\log|\gamma|.$ (45)
@@
$\lim_{n\to\infty}\mathbb{E}[|X_{n}|^{2}]$ $=\lim_{n\to\infty}\mathbb{E}[|V_{n}-\mathbb{E}[V_{n}|\widetilde{Y}_{1}^{n-1}]|^{2}]$ (47)
$=\lim_{n\to\infty}\frac{a_{n}^{2}+b_{n}^{2}+\|a_{n}D_{2,n-1}-b_{n}D_{1,n-1}\|^{2}}{|I_{2}+D_{n-1}^{T}D_{n-1}|}$
@@
where the third equality follows from the definitions of $a_{n}$ and $b_{n}$ together with (44) and (46).
@@
Together with (45) and (47), this immediately establishes that the desired results in (19) and (20) hold for Case B. The proof of the identity (48) is deferred to Appendix C. ∎
Now, combining Theorem 3.2 (the case $b=0$) and Theorem 4.1 (the case $b\neq 0$), we can characterize the maximum achievable rate of the SK(2) coding scheme, which in turn provides a lower bound on $C_{\mathrm{FB}}(P)$.
@@
$C_{\mathrm{SK2}}(P)=\max\left\{C_{\mathrm{SK1}}(P),\bar{I}_{\mathrm{SK2}}(P)\right\},$
where $C_{\mathrm{SK1}}(P)$ and $\bar{I}_{\mathrm{SK2}}(P)$ are given by (12) and (19), respectively.
@@
This spectral density corresponds to a stationary noise process of the form $Z_{i}=W_{i},\penalty 10000\ i\geq 1$, where $\{W_{i}\}_{i=1}^{\infty}$ is a zero-mean, unit-variance white Gaussian process. This corresponds precisely to the AWGN channel (4). It follows directly from Theorem 3.2 and Corollary 4.2 that for the case $b=0$, the SK(2) scheme reduces to the SK(1) scheme and is therefore optimal for this channel, as noted in Remark 3.3. For the case $b\neq 0$, the following corollary demonstrates that, for such channel, the optimal pair $(\gamma_{1},\gamma_{2})$ in the optimization problem (19)-(20) satisfies $|\gamma_{1}\gamma_{2}|=\sqrt{P+1}$. Consequently, the SK(2) coding scheme yields a broader family of optimal coding scheme for the AWGN channel.
@@
Consider the noise process $\{Z_{i}\}_{i=1}^{\infty}$ with the power spectral density given in (49). Then, the feedback capacity of the AWGN channel $Y_{i}=X_{i}+Z_{i},\penalty 10000\ i=1,2,$..., subject to the power constraint $P$, can be achieved via the SK(2) coding scheme defined in (15) and (14) for $b\neq 0$ with any parameter pair $(\gamma_{1},\gamma_{2})$ satisfying $|\gamma_{1}\gamma_{2}|=\sqrt{P+1}$.
@@
$\bar{I}_{\mathrm{SK2}}(\gamma_{1},\gamma_{2})\triangleq\log|\gamma_{1}\gamma_{2}|$
under the constraint (50).
Next, we consider the two cases. First, suppose the characteristic roots $(\gamma_{1},\gamma_{2})$ satisfy $\gamma_{1}=\overline{\gamma_{2}}$ and $|\gamma_{2}|=\gamma>1$. In this case, $\bar{I}_{\mathrm{SK2}}(\gamma_{1},\gamma_{2})=2\log|\gamma|$ and the power constraint (50) reduces to $\gamma^{4}-1\leq P$. Consequently,
@@
$\max_{\begin{subarray}{c}\gamma_{1},\gamma_{2}\in\mathbb{R}\\ \gamma_{1}\neq\gamma_{2}\end{subarray}}\bar{I}_{\mathrm{SK2}}(\gamma_{1},\gamma_{2})$ $=\max\log|\gamma_{1}\gamma_{2}|$
@@
where the maximum is attained by any pair $(\gamma_{1},\gamma_{2})$ satisfying $|\gamma_{1}\gamma_{2}|=\sqrt{P+1}$.
In summary, combining both cases we have shown that the maximum $\bar{I}_{\mathrm{SK2}}(P)$ in (19) is achieved whenever the characteristic roots $(\gamma_{1},\gamma_{2})$ satisfy $|\gamma_{1}\gamma_{2}|=\sqrt{P+1}$. In particular, the well-known feedback capacity of the AWGN channel, $\log(P+1)/2$, is attained by the SK(2) coding scheme for the case $b\neq 0$ with any parameter pair $(\gamma_{1},\gamma_{2})$ satisfying $|\gamma_{1}\gamma_{2}|=\sqrt{P+1}$, as desired. ∎
@@
with equality if and only if $\beta_{1}=0$. For different values of $\beta_{1}$ and $P=1,5$, we plot $C_{\mathrm{SK1}}(P)$, $\bar{I}_{\mathrm{SK2}}(P)$, and their difference as functions of $\beta_{1}$, as shown in Fig. 1.
Figure 1: Comparison of $C_{\mathrm{SK1}}(P)$ and $\bar{I}_{\mathrm{SK2}}(P)$ for stationary AR(1) Gaussian channels with $P=1,5$.
@@
$C_{\mathrm{SK2}}(P)=\bar{I}_{\mathrm{SK2}}(P)>C_{\mathrm{SK1}}(P).$
Thus, unlike for AR(1) Gaussian channels, the $\mathrm{SK}(1)$ coding scheme is not universally optimal. This disproves the conjecture of Butman [2] that in our notation $C_{\mathrm{SK1}}(P)=C_{\mathrm{FB}}(P)$.
Figure 2: Comparison of $C_{\mathrm{SK1}}(P)$ and $\bar{I}_{\mathrm{SK2}}(P)$ for stationary AR(2) Gaussian channels with $\beta_{2}=0.5,\penalty 10000\ 0.9$ and $P=1$.
@@
For any nonzero $x\in\mathbb{C}^{n}$, since $S$ is positive-definite Hermitian, we have
$x^{\ast}(S+UU^{\ast})x$ $=x^{\ast}Sx+x^{\ast}UU^{\ast}x$
$=x^{\ast}Sx+\|U^{\ast}x\|^{2}$
$\geq x^{\ast}Sx$